An Otto cycle with a compression ratio of 7.5 operates from the suction conditions of 14.2 psia, 85 o F.

Find the pressure and temperature at the end of compression, a) if cold air (k=1.4) is the working

substance, b) if hot air (k=1.32) is the working substance.


1
Expert's answer
2021-05-05T07:42:41-0400

given : "r_k =7.5"


"P_1 =14.2 Psi = 97.905 \frac{kN}{m^2}"


"T_1 = 85^0F = 302.6^0K"


if "k = 1.4"


"T_2 = r_k^{k-1} \times T_1= 7.5^{1.4-1}\times 302.6 = 677.47 ^0 K"


"P_2 =r_k \times \frac{T_2}{T_1}\times P_1 = 7.5 \times \frac{677.47}{302.6}\times 97.905 = 1643.94 \frac{kN}{m^2}"


if "k = 1.32"


"T_2 = r_k^{k-1} \times T_1= 7.5^{1.32-1}\times 302.6 = 576.62 ^0 K"


"P_2 =r_k \times \frac{T_2}{T_1}\times P_1 = 7.5 \times \frac{576.62}{302.6}\times 97.905 = 1399.22 \frac{kN}{m^2}"